# Early stages of collective cell invasion: Biomechanics

### v1

@@ 4. Discussion
We have proposed a biomechanical model of early invasion of cells that come from aggregates. Its basis is the cellular Potts model that incorporates durotaxis vanOers modified to include active forces that affect differently mesenchymal (M), epithelial (E) and hybrid E/M cells. As we are interested in early stages, we have not included cellular proliferation. We have not included biochemical drivers such as chemotaxis, cellular Notch signaling, cancer stem cells or the EMT boc18 , which are left for a subsequent publication. Once the EMT model and the most relevant biochemical descriptions are determined for the particular cancer to study boc18 ; vil22 ; muk22 ; hir23 , these biochemical processes determine cell phenotypes that are assigned different mechanical characteristics as discussed in the present paper. This paper has implications for metastasis but our predictions are qualitative rather than quantitative until they are supplemented by future work.
@@
As in previous works vanOers ; veg20 , we have assumed that cells act as contractile units resulting in a first moment of area representation for their force distribution lem10 . The resulting traction forces of the cells are realistic (but they are not the gradients of the CPM Hamiltonian) and the strains in the ECM are calculated via the finite-element method vanOers . This approach produces cells moving by durotaxis. At a higher computational cost, traction forces consistent with the variations $\Delta H$ of the Hamiltonian in the CPM can be computed from the gradients of $H$ using finite differences, interpolations and smooting of pixels in the CPM grid ren19 . Here have split each time step into two: For half a step, we have adopted the simpler durotaxis description of Ref. vanOers and for the other half step we have used a simple model of active forces. Including more realistic active migration forces (e.g., nonlinear push and/or pull forces) is possible. For example, if certain cells move by chemotaxis in the gradient of a continuum field as in angiogenesis veg20 , the gradient of the corresponding term in the CPM Hamiltonian (calculated using finite differences, interpolation and smoothing ren19 ) produces active chemotactic forces. These forces could be incorporated via one half of a fractional time step while the traction forces of the cells comprise the other half step as in the present paper.
@@
Supporting information
Video S1.
Active and passive single cells.
Video S2.
Migration force in active cells.
Video S3.
Migration patterns.
Video S4.
Adhesive, active and hybrid cells.
Video S5.
Collective invasion of E cells from a fingering instability.
Video S6.
Collective invasion of E cells from a finger led by a basal cell.
@@ Acknowledgments
This work has been supported by the FEDER/Ministerio de Ciencia, Innovación y Universidades – Agencia Estatal de Investigación (MCIN/ AEI/10.13039/501100011033) grants PID2020-112796RB-C22 and PID2024-155528OB-C22. Funding for APC: Universidad Carlos III de Madrid (Agreement CRUE-Madroño 2025).
Appendix A Flowcharts of the fractional step and single step methods
Start: Increment nnCompute cell displacement due to ECM stiffness (FEM with force $f$)Compute $\Delta H_{strain}$ using displacement and deformation tensor CPM moves: $\Delta H=\Delta H_{volume}+\Delta H_{contact}+\Delta H_{strain}$; accept/reject via Monte Carlo End of first fractional sub-stepCompute cell displacement due to ECM migration force $f_{m}$ (FEM)Compute $\Delta H_{migration}$ using displacement and deformation tensor CPM moves: $\Delta H=\Delta H_{volume}+\Delta H_{contact}+\Delta H_{migration}$; accept/reject via Monte Carlo Update cell pixel identities (phenotype assignment)End of fractional iteration / ready for next increment
Start: Increment nnCompute cell displacement due to total ECM forces $f+f_{m}$ (FEM)Compute $\Delta H_{strain+migration}$ using displacement and deformation tensor CPM moves: $\Delta H=\Delta H_{volume}+\Delta H_{contact}+\Delta H_{strain+migration}$; accept/reject via Monte Carlo Update cell pixel identities (phenotype assignment)End of single step iteration / ready for next increment

### v2

@@ VI Discussion
We have proposed a biomechanical model of early invasion of cells issuing from aggregates. Its basis is the cellular Potts model incorporating durotaxis vanOers , modified to include active migration forces that affect differently mesenchymal M, epithelial E, and hybrid E/M cells. We are interested in early stages before cellular proliferation. Biochemical drivers such as chemotaxis, cellular Notch signaling, cancer stem cells or the EMT boc18 , will be left for a subsequent publication. Once the EMT model and the most relevant biochemical descriptions are determined for the particular cancer to study boc18 ; vil22 ; muk22 ; hir23 , they determine cell phenotypes that have different mechanical characteristics. Our predictions have implications for metastasis but they are qualitative rather than quantitative until supplemented by future work.
@@
As in previous works vanOers ; veg20 , we have assumed that cells act as contractile units resulting in a first moment of area representation for their force distribution lem10 . The resulting traction forces of the cells are realistic (but they are not the gradients of the CPM Hamiltonian), the strains in the ECM are calculated via the finite-element method and modify the substrate stiffness while keeping the extension-retraction symmetry vanOers . This approach produces cells moving by durotaxis. Adding active forces to traction forces interferes with the extension-retraction symmetry and may impede cell migration in some cases. Here have split each MCTS into two: For half a step, we have adopted the simpler durotaxis description of Ref. vanOers and for the other half step we have used a simple active forces that direct cells towards an attraction center. Including more realistic active migration forces (e.g., nonlinear push and/or pull forces) is possible. For example, if certain cells move by chemotaxis in the gradient of a continuum field as in angiogenesis veg20 , the gradient of the corresponding term in the CPM Hamiltonian (calculated using finite differences, interpolation and smoothing ren19 ) produces active chemotactic forces. These forces could be incorporated via one half of a fractional time step while the traction forces of the cells comprise the other half step, as in the present paper.
@@ Acknowledgments
This work has been supported by the FEDER/Ministerio de Ciencia, Innovación y Universidades – Agencia Estatal de Investigación (MCIN/AEI/10.13039/501100011033) grants PID2020-112796RB-C22 and PID2024-155528OB-C22.
Appendix A Parameters and dynamic solutions
Here we enumerate 21 configurations and a table with parameter values corresponding to these configurations that produce seven different patterns depicted in Fig. 2.
@@ Appendix A Parameters and dynamic solutions
Appendix B Flowcharts of the fractional step and single step methods
Start: Increment nnCompute cell displacement due to ECM stiffness (FEM with force $f$)Compute $\Delta H_{\text{duro}}$ using displacement and deformation tensor CPM moves: $\Delta H=\Delta H_{\text{volume}}+\Delta H_{\text{contact}}+\Delta H_{\text{duro}}$; accept/reject via Monte Carlo End of first fractional sub-stepCompute cell displacement due to ECM migration force $f_{m}$ (FEM)Compute $\Delta H_{\text{mig}}$ using displacement and deformation tensor CPM moves: $\Delta H=\Delta H_{\text{volume}}+\Delta H_{\text{contact}}+\Delta H_{\text{mig}}$; accept/reject via Monte Carlo Update cell pixel identities (phenotype assignment)End of fractional iteration / ready for next increment
Start: Increment nnCompute cell displacement due to total ECM forces $f+f_{m}$ (FEM)Compute $\Delta H_{\text{duro+mig}}$ using displacement and deformation tensor CPM moves: $\Delta H=\Delta H_{\text{volume}}+\Delta H_{\text{contact}}+\Delta H_{\text{duro+mig}}$; accept/reject via Monte Carlo Update cell pixel identities (phenotype assignment)End of single step iteration / ready for next increment